# Looping Around Neutrino Charge Radius at Ultra-Near Reactor Experiments

### v1

@@ Introduction
Elements of the leptonic mixing matrix and neutrino mass squared differences are not the only parameters that can be probed at reactor experiments. Specifically, electromagnetic neutrino properties can also be studied [7, 8]. Since neutrinos are not charged particles, there are no electromagnetic interactions involving neutrinos at tree level; however, this changes at the quantum level, where magnetic, electric, and anapole moments appear, along with the neutrino charge radius. The first two are proportional to neutrino mass and hence negligible in the Standard Model (SM) framework, while the anapole moment and neutrino charge radius do not vanish for and hence are appealing targets for being tested at neutrino experiments. In the massless neutrino limit, the anapole moment is equal to the neutrino charge radius, up to an factor [9]. The strongest reported reactor neutrino constraints on the neutrino charge radius come from the Krasnoyarsk [10] and TEXONO [11] experiments; these limits are set on the electron neutrino charge radius and are only about an order of magnitude larger than the SM value. This motivates the question whether near-future reactor neutrino experiments could discover the neutrino charge radius through – scattering, which will be addressed in this work. While we will focus exclusively on reactor experiments, it is worthwhile to acknowledge recent neutrino charge radius studies in connection to other types of experiments [12, 13, 14, 15], some of which focus on the muon neutrino charge radius.
In addition to radiative corrections associated with neutrino charge radius, we will consider further radiative corrections in – scattering, presented in detail in Refs. [16, 17]; for other studies on probing next-to-leading-order effects at neutrino experiments, see [18, 19, 18, 20]. We will, in fact, demonstrate that these additional radiative corrections are highly relevant and impact our results both qualitatively and quantitatively. Regarding the former, these contributions cannot be isolated from the charge radius at reactor neutrino experiments, complicating the extraction of the neutrino charge radius. Regarding the latter, they counteract the charge radius contribution to the – cross section, making discovery at reactor experiments more challenging.
@@ II.2 Neutrino-electron Scattering
Equation (7) is only for tree-level calculations. Throughout this paper, we refer to it as the leading-order (LO) cross section. The radiative correction due to the CR can be readily included by adding the following correction to the LO differential cross section [16, 17]
where is the fine-structure constant. This effect is equivalent to a shift of in the LO cross section,
which is the approach often used in the literature [27, 31, 12, 32].
@@ II.3 Radiative Corrections
The full calculation at one-loop order, as done in Refs. [17, 16] using the effective field theory, yields the following values of the coefficients and (at renormalization scale GeV) for electron antineutrinos scattering off electrons
The other input parameters such as , , , and , are also determined at GeV scale.333Numerical values can be found in Refs. [17, 16]. Given that, we are armed with all input parameters necessary for calculating the cross section featuring both tree-level and radiative contributions.
@@ IV.1 Reactor Fluxes
To generate the reactor fluxes, we follow Ref. [42], and construct the antineutrino spectra (for 235U, 238U, 239Pu, and 241Pu) from a order polynomial parametrization [43]. We assume a consistent load factor of 80%, and consider the fuel fractions shown in Table 1. In our analysis we will assume reactor flux uncertainties of 0.5% and 1%, respectively. While in the last decade there were several predictions that differed by up to %, more recent flux calculations have led to reactor antineutrino anomaly fading away [44, 45]. In addition, data-driven methods for flux determination led by PROSPECT [46], and which will be joined in the future by both TAO and CLOUD, will improve on our knowledge of the flux even further.
@@ IV.2 Detector Configurations and Backgrounds
Background considerations are important in order to assess the realistic capability of TAO and CLOUD for probing NLO effects. In our analysis, we chiefly follow Ref. [40] where it was shown that the main backgrounds come from natural radioactivity and cosmic muons which produce spallation isotopes (i.e. cosmogenic background). In Ref. [40], the latter dominates over the former for a typical liquid scintillator detector with a 300 mwe overburden. In our work, since both detectors have significantly less overburden, the cosmogenic background is higher than in the scenario considered in Ref. [40]. Fortunately, the detector-reactor distance is also much shorter in our work compared to Ref. [40]. The resulting higher neutrino flux compensates the loss caused by the cosmogenic background. This also implies that in our work, the backgrounds due to natural radioactivity would be more subdominant.
For the TAO detector, we conservatively estimate that the shielding corresponds to mwe. In order to take that into account, we adopt cosmogenic backgrounds at 300 mwe from Ref. [40], and scale them to mwe using Table VII in the same work.
@@ Acknowledgements.
- [42] M. Baldoncini, I. Callegari, G. Fiorentini, F. Mantovani, B. Ricci, V. Strati, and G. Xhixha, Reference worldwide model for antineutrinos from reactors, Physical Review D 91 (Mar., 2015) 065002. arXiv:1411.6475 [hep-ex, physics:physics].
- [43] T. A. Mueller, D. Lhuillier, M. Fallot, A. Letourneau, S. Cormon, M. Fechner, L. Giot, T. Lasserre, J. Martino, G. Mention, A. Porta, and F. Yermia, Improved predictions of reactor antineutrino spectra, Physical Review C 83 (May, 2011) 054615.

### v2

@@ Introduction
Elements of the leptonic mixing matrix and neutrino mass squared differences are not the only parameters that can be probed at reactor experiments. Specifically, electromagnetic neutrino properties can also be studied [7, 8]. Since neutrinos are not charged particles, there are no electromagnetic interactions involving neutrinos at tree level; however, this changes at the quantum level, where magnetic, electric, and anapole moments appear, along with the neutrino charge radius. The first two are proportional to neutrino mass and hence negligible in the Standard Model (SM) framework, while the anapole moment and neutrino charge radius do not vanish for and hence are appealing targets for being tested at neutrino experiments. In the massless neutrino limit, the anapole moment is equal to the neutrino charge radius, up to an factor [9]. The strongest reported reactor neutrino constraints on the neutrino charge radius come from the Krasnoyarsk [10] and TEXONO [11] experiments; these limits are set on the electron neutrino charge radius and are only about an order of magnitude larger than the SM value. This motivates the question whether near-future reactor neutrino experiments could discover the neutrino charge radius through – scattering, which will be addressed in this work. While we will focus exclusively on reactor experiments, it is worthwhile to acknowledge recent neutrino charge radius studies in connection to other types of experiments [12, 13, 14, 15, 16], some of which focus on the muon neutrino charge radius.
In addition to radiative corrections associated with neutrino charge radius, we will consider further radiative corrections in – scattering, presented in detail in [17, 18]; for other studies on probing next-to-leading-order effects at neutrino experiments, see [19, 20, 19, 21]. We will, in fact, demonstrate that these additional radiative corrections are highly relevant and impact our results both qualitatively and quantitatively. Regarding the qualitative impact, these contributions cannot be isolated at reactor neutrino experiments, complicating the extraction of the neutrino charge radius. Regarding the quantitative impact, they counteract the charge radius contribution to the – cross section, making discovery at reactor experiments more challenging.
@@ II.2 Neutrino-electron Scattering
Equation (7) is only for tree-level calculations. Throughout this paper, we refer to it as the leading-order (LO) cross section. At the next-to-leading order (NLO), there are radiative corrections, with the CR being the most prominent one. The CR correction is often included via a shift of in the literature [28, 32, 12, 33]:
This shift is equivalent to the following correction to the LO differential cross section [17, 18]:
where is the fine-structure constant. Note that Eq. (10) is an correction. The shift of could also generate an correction which is neglected here for a consistent truncation of radiative corrections. We have numerically checked that neglecting the correction causes vanishingly small variation of our results.
@@ II.3 Radiative Corrections
The full calculation at one-loop order, as done in [18, 17] using the effective field theory, yields the following values of the coefficients and (at renormalization scale GeV) for electron antineutrinos scattering off electrons
The other input parameters such as , , , and , are also determined at GeV scale.333Numerical values can be found in [18, 17]. Given that, we are armed with all input parameters necessary for calculating the cross section featuring both tree-level and radiative contributions.
@@ IV.1 Reactor Fluxes
To generate the reactor fluxes, we follow [43], and construct the antineutrino spectra (for 235U, 238U, 239Pu, and 241Pu) from a order polynomial parametrization [44]. We assume a consistent load factor of 80%, and consider the fuel fractions shown in Table 1. For a review and comparison of reactor antineutrino spectra, see [45]. In our analysis we will assume reactor flux uncertainties of 0.5% and 1% [46], respectively. While in the last decade there were several predictions that differed by up to %, more recent flux calculations have led to reactor antineutrino anomaly fading away [47, 48]. However, to reach the reactor flux uncertainties of 0.5% and 1%, we will rely on data-driven methods for flux determination led by PROSPECT [49], and joined by both TAO and CLOUD, which will measure the fluxes in-situ and further improve our knowledge of the reactor fluxes.
@@ IV.2 Detector Configurations and Backgrounds
Background considerations are important in order to assess the realistic capability of TAO and CLOUD for probing NLO effects. In our analysis, we chiefly follow [41] where it was shown that the main backgrounds come from natural radioactivity and cosmic muons which produce spallation isotopes (i.e. cosmogenic background). In [41], the latter dominates over the former for a typical liquid scintillator detector with a 300 mwe overburden. In our work, since both detectors have significantly less overburden, the cosmogenic background is higher than in the scenario considered in [41]. Fortunately, the detector-reactor distance is also much shorter in our work compared to [41]. The resulting higher neutrino flux compensates the loss caused by the cosmogenic background. This also implies that in our work, the backgrounds due to natural radioactivity would be more subdominant.
For the TAO detector, we conservatively estimate that the shielding corresponds to mwe. In order to take that into account, we adopt cosmogenic backgrounds at 300 mwe from [41], and scale them to mwe using Table VII in the same work.
@@ Acknowledgements.
- [15] S.-F. Ge and P. Pasquini, Disentangle neutrino electromagnetic properties with atomic radiative pair emission, JHEP 12 (2023) 083, [2306.12953].
@@
- [43] M. Baldoncini, I. Callegari, G. Fiorentini, F. Mantovani, B. Ricci, V. Strati, and G. Xhixha, Reference worldwide model for antineutrinos from reactors, Physical Review D 91 (Mar., 2015) 065002, [1411.6475].
- [44] T. A. Mueller, D. Lhuillier, M. Fallot, A. Letourneau, S. Cormon, M. Fechner, L. Giot, T. Lasserre, J. Martino, G. Mention, A. Porta, and F. Yermia, Improved predictions of reactor antineutrino spectra, Physical Review C 83 (May, 2011) 054615, [1101.2663].
- [45] L. Périssé, A. Onillon, X. Mougeot, M. Vivier, T. Lasserre, A. Letourneau, D. Lhuillier, and G. Mention, Comprehensive revision of the summation method for the prediction of reactor fluxes and spectra, Phys. Rev. C 108 (2023) 055501, [2304.14992].
- [46] “JUNO-TAO Status and Prospect.” https://indico.cern.ch/event/1291157/contributions/5899076/attachments/2899988/5085319/TAO_Prague_240719-7.pdf.